# Comprehensive X-ray Observations of the Exceptional Ultra-long X-ray and Gamma-ray Transient GRB 250702B with Swift, NuSTAR, and Chandra: Insights from the X-ray Afterglow Properties

### v2

@@ Abstract
journal: ApJLfacilities: Neil Gehrels Swift Observatory, Nuclear Spectroscopic Telescope Array, Chandra X-ray Observatory software: Astropy (Astropy Collaboration et al., 2018, 2022), SciPy (scipy), emcee (D. Foreman-Mackey et al., 2013), corner (D. Foreman-Mackey, 2016), stingray (D. Huppenkothen et al., 2019), XSPEC (K. A. Arnaud, 1996), GUANO (Tohuvavohu2020), NITRATES (J. DeLaunay & A. Tohuvavohu, 2022), BatAnalysis (batanalysis), HEASoft, CIAO
@@ 1 Introduction
An alternative origin for some ultra-long events is that they arise from relativistic jetted tidal disruption events (TDEs) in which a star is shredded by a massive black hole, launching a jet along our line of sight (J. S. Bloom et al., 2011; Levan2011; S. B. Cenko et al., 2012; Pasham2015; I. Andreoni et al., 2022; Pasham2023). This was the case for GRB 110328A/Sw J1644+57 (D. N. Burrows et al., 2011; J. S. Bloom et al., 2011), which was the first event discovered in this class. Unlike classical GRBs, the recognized sample of relativistic TDEs lacks highly variable prompt gamma-ray emission, and these events were instead discovered via longer-duration “image” triggers (J. R. Cummings et al., 2011; Sakamoto2011J1644) or searches using multi-day stacked datasets (S. B. Cenko et al., 2012; G. C. Brown et al., 2015).
@@ 2.2 NuSTAR
We used standard tasks within the NuSTAR Data Analysis Software pipeline (NuSTARDAS; CALDB 20250714) within HEASoft to reprocess the data, and then extract lightcurves and spectra. In the first epoch, the X-ray counterpart is clearly detected with a mean count rate of $\sim$ 0.3 cts s-1 per FPM (nustargcn). We extracted the source counts from a circular region with radius 45″ from both FPMA and FPMB, and utilized a source free region with radius 120″ to extract the background. The background region was placed on the same detector as the transient. After subtracting the background, the mean count rate is $\sim$ 0.25 cts s-1 per FPM.
@@
Figure 1: Left: X-ray lightcurve of GRB 250702B combining Swift and Chandra in the 0.30.3-$10$ keV energy range. The start time $T_{0}$ is the GBM “D” burst. The red lines show the best-fit decay from the initial trigger time of GRB 250702D (Table 2). The dashed red line ($t^{-1.9}$) includes the first XRT orbit in the fit, whereas the solid red line excludes the first XRT orbit and provides a better fit to the late-time data ($t^{-1.8}$). The observation windows of our NuSTAR data are shown as orange shaded regions. For reference we show the expected slopes for a relativistic TDE, corresponding to both complete ($t^{-5/3}$) and partial ($t^{-9/4}$) disruption. The bottom panel shows the 0.30.3-$10$ keV lightcurve fit residuals relative to the solid red line. Right: Histogram showing the posterior probability density of the best-fit onset time for the X-ray emission. The gray histogram shows the $T_{0}$ posterior obtained when fitting all XRT data, and the red histogram shows the $T_{0}$ posterior when excluding the first XRT orbit. The trigger times reported by Fermi (red) and EP (blue) are shown as vertical lines (Table 2). The July 1 window in which EP reports an initial detection is to the left of the vertical blue line. We also show as vertical black lines the times of Swift/BAT gamma-ray detections. All times are relative to the trigger of GRB 250702D at MJD 60858.548.
@@ 2.3 Chandra X-ray Observatory
We carried out a late-time X-ray observation with the Chandra X-ray Observatory (CXO) through a DDT request (PI: O’Connor; ObsID: 31011) starting on 2025-08-09 at 05:49:28 UT (corresponding to $\sim$ 37.69 d after discovery) for $\sim$$27.7$ ks using ACIS-S (chandragcn). The target was centered at the aimpoint of the S3 chip. The Chandra data were retrieved from the Chandra Data Archive (CDA) [^4] https://cda.harvard.edu/chaser/. We re-processed the data using the CIAO v4.17.0 data reduction package with CALDB v4.11.6. At the subarcsec location of the near-infrared counterpart (Levan2025), we identify a clear X-ray source. We find 25 counts (0.50.5-88 keV) in a circular source region with radius $1.5\arcsec$ centered on the localization obtained with wavdetect. We extracted background counts from a nearby source free region with radius $30\arcsec$. Using the best fit X-ray spectrum (§3.2), we derive an unabsorbed 0.30.3-$10$ keV flux of $(3.3^{+0.7}_{-0.6})\times 10^{-14}$ erg cm-2 s-1. The Chandra data confirms that there are no contaminating point sources within either the Swift or NuSTAR source extraction regions.
Another DDT observation (PI: Eyles-Ferris; R. A. J. Eyles-Ferris et al. 2025; R. A. Eyles-Ferris et al. 2025) occurred starting on 2025-09-05 at 23:26:41 UT, corresponding to 65.5 d after discovery, using ACIS-S for a total of $39.55$ ks. We analyzed this observation in the same manner, finding 12 counts in a $1.5\arcsec$ radius, and obtain an unabsorbed 0.30.3-$10$ keV flux of $(1.15^{+0.40}_{-0.30})\times 10^{-14}$ erg cm-2 s-1.
@@ 3.1.1 Temporal Modeling
We modeled the 0.30.3-$10$ keV X-ray lightcurve (Swift/XRT and Chandra) with a powerlaw of the form $t^{-\alpha}$ using the emcee package (D. Foreman-Mackey et al., 2013). As the exact onset time $T_{0}$ of the transient is unclear (Table 2), we performed multiple different fits to the data to explore the range of possible onset times and temporal decay rates. We first allowed the onset time to be a free parameter, and then performed an additional fit with the onset time fixed to the first high-energy trigger time of GRB 250702D by Fermi (see Table 2).
@@
We performed these fits both with and without the inclusion of the initial XRT orbit at <<$0.55$ days after discovery (see Figure 14). We demonstrate that the initial XRT observation shows flaring activity (§3.1.2) and may bias the temporal analysis. Ignoring the initial flaring episode does not significantly change the inferred decay slope (see Figure 14), but does shift the inferred onset time (Figure 1; right panel) towards earlier times. This demonstrates that the exact decay slope and onset time depend strongly on the assumptions taken for the origin of the X-ray lightcurve.
@@
Using the trigger of GRB 250702D as the onset time, we similarly modeled the NuSTAR 33-$79$ keV flux (see §3.2) across the three epochs (with mid-times of 1.58, 5.70, 9.98 d; Table 1) with a temporal powerlaw model. We derive a temporal index of $\alpha$$2.26\pm 0.15$. This is slightly steeper, but consistent at the $\sim$$2\sigma$ level, with the XRT fit using the same fixed start time (see Figure 14). We note that the NuSTAR data is limited to three epochs and the third epoch suffers from large error bars, which suggests that the slope derived from the combined Swift and Chandra lightcurve is more accurate. The first epoch of NuSTAR data also suffers from flaring (see §3.1.2), which can also explain its higher flux that can lead to a steeper inferred slope. The shallower slope measured from Swift and Chandra is also more consistent with the early Swift/BAT survey mode upper limits (see §2.1.1; Table 3). This is shown in Figure 5. While the NuSTAR extrapolation goes through two BAT detections of the prompt emission, this depends strongly on the assumed temporal slope and we emphasize the BAT detections are not included in the fit.
Figure 5: Left: Swift/BAT survey mode (1414-$195$ keV) on emission from GRB 250702B from the start of July 1, 2025 to the end of July 4, 2025. Black points show the Swift/BAT detections (1414-$195$ keV; Tables 5 and 4) from both GUANO and survey data. Gray downward triangles represent $5\sigma$ upper limits from Swift/BAT in the 1414-$195$ keV energy range. The trigger times reported by Fermi (red) and EP (blue) are shown as vertical lines (Table 2). The July 1 window in which EP reports an initial detection is shown as a blue shaded region. The time window of our initial NuSTAR observation (starting 1.3 d after the initial Fermi trigger) is shown as an orange shaded region. Right: Best-fit temporal powerlaw to the NuSTAR lightcurve in 33-$79$ keV. The $1\sigma$ error on the fit is shown as a shaded orange region. Swift/BAT upper limits are shown as gray triangles and detections as black circles. The start time is taken as the trigger time of GRB 250702D. The vertical red lines show the Fermi trigger times of GRBs 250702B, 250702C, and 250702E (see Table 2).
@@ 3.2.1 X-ray Spectra
In all fits we have frozen the Galactic hydrogen column density to $N_{\textrm{H,gal}}$$3.34\times 10^{21}$ cm-2 (Willingale2013). The data were modeled between 0.3100.3-10 keV for Swift and 3793-79 keV for NuSTAR. Using an absorbed powerlaw model, we find an intrinsic hydrogen column density $N_{\textrm{H,z}}$$(6.2\pm 0.8)\times 10^{22}$ cm-2 and X-ray photon index $\Gamma$$1.82\pm 0.03$ with Cstat/dof ==$1277/1390$. The spectral fit and fit residuals are shown in Figure 7. We do not find evidence for spectral features, such as emission or absorption lines or reflection features. This further supports an extragalactic scenario for GRB 250702B as it would be peculiar for an X-ray binary in outburst.
@@ 4 The Peculiar High-Energy Properties of GRB 250702B
Figure 10: K-corrected rest frame (0.30.3-$10$ keV) X-ray lightcurves of Swift long duration GRBs, ultra-long GRBs (GRBs 111209A, 121027A, 130925A; Levan2014; P. A. Evans et al., 2014a), and relativistic jetted TDEs (J. S. Bloom et al., 2011; S. B. Cenko et al., 2012; G. C. Brown et al., 2015; Pasham2023) compared to GRB 250702B for a few onset times. For the July 2nd onset time (purple), we have used the trigger time of GRB 250702D as $T_{0}$ (Table 2). For the July 1 onset time (blue), we use the beginning of the first EP detection window (Hui Sun, private communication).
@@ 4.1 Ultra-long Prompt Emission
While these properties are atypical for GRBs, they have likewise never been observed from a relativistic TDE. While relativistic TDEs have been uncovered in binned gamma-ray data (S. B. Cenko et al., 2012; G. C. Brown et al., 2015), or 1000 s long image triggers (Sakamoto2011J1644), they have never triggered high-energy monitors in the standard “rate” trigger method that detects typical GRBs, and they have definitely not done so multiple times in a 24 hour window. As near all-sky gamma-ray monitors have been operating for decades, multiple gamma-ray triggers would be more difficult to miss in the past than early soft X-ray emission that could only have been found by the Einstein Probe which launched only in the last 2 years (Yuan2025). A detailed sub-threshold gamma-ray search in the day(s) preceding ultra-long GRBs is likely necessary to further investigate this possibility.
@@
An important clue can be obtained from the isotropic-equivalent gamma-ray energy $E_{\gamma,\textrm{iso}}$. The integrated energy obtained by Konus-Wind (D. Frederiks et al., 2025) is $E_{\gamma,\textrm{iso}}$$(1.4^{+0.4}_{-0.2})\times 10^{54}$ erg in the 11-$10,000$ keV energy range (see also Neights2025). This is at the high end of the isotropic-equivalent energies found in classical GRBs and ultra-long GRBs (see Figure 11). This is also substantially higher than the energy released ($\approx$$10^{51}$ erg; 11-$10,000$ keV) by Sw J1644+57 during its initial Swift/BAT detection (Sakamoto2011J1644), which is shown as an orange shaded region in Figure 11. It is possible Sw J1644+57 released more (weaker) gamma-ray pulses that went undetected, so its inferred $E_{\gamma}$ is a lower limit. However, the gamma-ray emission from Sw J1644+57 was substantially softer (peak energy $\sim$$70$ keV; Levan2011) than the >>1 MeV emission observed from GRB 250702B. Moreover, the individual energetics of the “DBE” events yield values a factor of 10 higher than observed from Sw J1644+57.
@@
In addition, the Fermi data revealed a critical piece of evidence in the minimum variability time. Neights2025 identified a very short second duration ($\lesssim$11 s; observer frame) variability timescale, which is a factor of 100100-$1,000$ smaller than observed in relativistic TDEs (J. S. Bloom et al., 2011; Pasham2023). Relativistic TDEs, and accreting black holes in general, are limited by the Schwarzschild light crossing time $t_{s}$$2R_{\textrm{g}}/c$, where $R_{\textrm{g}}$$GM/c^{2}$ is the gravitational radius, which depends on the black hole mass. As larger masses lead to longer predicted minimum variability timescales, if a massive black hole is invoked it would require a small $\lesssim 5\times 10^{4}M_{\odot}$ intermediate-mass black hole. It is important to note that the light crossing timescale argument is only a lower limit to the minimum variability timescale, and the black hole size determined through this argument for past relativistic TDEs is larger than the latest analysis based on their jet shutoff timescales (T. Eftekhari et al., 2024).
@@
However, this seconds long variability is typical of that observed in long duration GRBs, and would therefore most obviously favor an interpretation that invokes a stellar-mass progenitor such as in GRBs. As an additional argument, in the collapsar progenitor channel, the central engine for long GRBs is likely a stellar-mass black hole (Woosley1993). Assuming a $\sim 5M_{\odot}$ black hole, the predicted minimum variability timescale based on the Schwarzschild light crossing time is $5\times 10^{-5}$ s, which is significantly smaller (by orders of magnitude) than the minimum variability observed in GRBs (see, e.g., $^{-2}-1001$$0$ s; V. Z. Golkhou & N. R. Butler, 2014). Following this argument, if applied to massive black holes, it would imply the minimum variability from light crossing is significantly smaller than the likely true intrinsic variability that can be observed. We therefore consider the second duration minimum variability of GRB 250702B (Neights2025) as strong evidence for a stellar-mass progenitor (§5.2).

### v3

@@ Abstract
journal: ApJLfacilities: Neil Gehrels Swift Observatory, Nuclear Spectroscopic Telescope Array, Chandra X-ray Observatory software: Astropy (Astropy Collaboration et al., 2018, 2022), SciPy (scipy), emcee (D. Foreman-Mackey et al., 2013), corner (D. Foreman-Mackey, 2016), stingray (D. Huppenkothen et al., 2019), XSPEC (K. A. Arnaud, 1996), GUANO (Tohuvavohu2020), NITRATES (J. DeLaunay & A. Tohuvavohu, 2022), BatAnalysis (batanalysis), CIAO (A. Fruscione et al., 2006), HEASoft (2014ascl.soft08004N)
@@ 1 Introduction
An alternative origin for some ultra-long events is that they arise from relativistic jetted tidal disruption events (TDEs) in which a star is shredded by a massive black hole, launching a jet along our line of sight (J. S. Bloom et al., 2011; Levan2011; S. B. Cenko et al., 2012; Pasham2015; I. Andreoni et al., 2022; Pasham2023). This was the case for GRB 110328A/Sw J1644+57 (D. N. Burrows et al., 2011; J. S. Bloom et al., 2011), which was the first event discovered in this class. Unlike classical GRBs, the recognized sample of relativistic TDEs lacks highly variable prompt gamma-ray emission, and these events were instead discovered via longer-duration “image” triggers (J. R. Cummings et al., 2011; Sakamoto2011J1644; D. N. Burrows et al., 2011) or searches using multi-day stacked datasets (S. B. Cenko et al., 2012; G. C. Brown et al., 2015).
@@ 2.2 NuSTAR
We used standard tasks within the NuSTAR Data Analysis Software pipeline (NuSTARDAS; CALDB 20250714) within HEASoft (2014ascl.soft08004N) to reprocess the data, and then extract lightcurves and spectra. In the first epoch, the X-ray counterpart is clearly detected with a mean count rate of $\sim$ 0.3 cts s-1 per FPM (nustargcn). We extracted the source counts from a circular region with radius 45″ from both FPMA and FPMB, and utilized a source free region with radius 120″ to extract the background. The background region was placed on the same detector as the transient. After subtracting the background, the mean count rate is $\sim$ 0.25 cts s-1 per FPM.
@@
Figure 1: Left: X-ray lightcurve of GRB 250702B combining Swift and Chandra in the 0.30.3-$10$ keV energy range. The start time $T_{0}$ is the GBM “D” burst. The red lines show the best-fit decay from the initial trigger time of GRB 250702D (Table 2). The dashed red line ($t^{-1.9}$) includes the first XRT orbit in the fit, whereas the solid red line excludes the first XRT orbit and provides a better fit to the late-time data ($t^{-1.8}$). The observation windows of our NuSTAR data are shown as orange shaded regions. For reference we show the expected slopes for a relativistic TDE, corresponding to both complete ($t^{-5/3}$) and partial ($t^{-9/4}$) disruption. The bottom panel shows the 0.30.3-$10$ keV lightcurve fit residuals relative to the solid red line. Right: Histogram showing the posterior probability density of the best-fit $T_{0}$ time for the X-ray afterglow emission. The gray histogram shows the $T_{0}$ posterior obtained when fitting all XRT data, and the red histogram shows the $T_{0}$ posterior when excluding the first XRT orbit. The trigger times reported by Fermi (red) and EP (blue) are shown as vertical lines (Table 2). The July 1 window in which EP reports an initial detection is to the left of the vertical blue line. We also show as vertical black lines the times of Swift/BAT gamma-ray detections. All times are relative to the trigger of GRB 250702D at MJD 60858.548.
@@ 2.3 Chandra X-ray Observatory
We carried out a late-time X-ray observation with the Chandra X-ray Observatory (CXO) through a DDT request (PI: O’Connor; ObsID: 31011) starting on 2025-08-09 at 05:49:28 UT (corresponding to $\sim$ 37.69 d after discovery) for $\sim$$27.7$ ks using ACIS-S (chandragcn). The target was centered at the aimpoint of the S3 chip. The Chandra data were retrieved from the Chandra Data Archive (CDA) [^4] https://cda.harvard.edu/chaser/. We re-processed the data using the CIAO v4.17.0 data reduction package (A. Fruscione et al., 2006) with CALDB v4.11.6. At the subarcsec location of the near-infrared counterpart (Levan2025), we identify a clear X-ray source. We find 25 counts (0.50.5-88 keV) in a circular source region with radius $1.5\arcsec$ centered on the localization obtained with wavdetect. We extracted background counts from a nearby source free region with radius $30\arcsec$. Using the best fit X-ray spectrum (§3.2), we derive an unabsorbed 0.30.3-$10$ keV flux of $(3.3^{+0.7}_{-0.6})\times 10^{-14}$ erg cm-2 s-1. The Chandra data confirms that there are no contaminating point sources within either the Swift or NuSTAR source extraction regions.
Additional Chandra DDT observations were obtained and we re-analyze them here. We analyzed these observation in the same manner outlined above. The first (PI: Li; EP250702a-arxiv) began on 2025-07-18 at 19:24:33 UT, corresponding to 16.8 d after discovery, using ACIS-I for 14.9 ks. We find 40 counts in a $1.5\arcsec$ radius, and obtain an unabsorbed 0.30.3-$10$ keV flux of $(1.2\pm 0.2)\times 10^{-13}$ erg cm-2 s-1 assuming out best fit X-ray spectrum (§3.2). The third DDT observation (PI: Eyles-Ferris; R. A. J. Eyles-Ferris et al. 2025; R. A. Eyles-Ferris et al. 2025) started on 2025-09-05 at 23:26:41 UT, corresponding to 65.5 d after discovery, using ACIS-S for a total of $39.55$ ks. We find 12 counts in a $1.5\arcsec$ radius, and obtain an unabsorbed 0.30.3-$10$ keV flux of $(1.15^{+0.40}_{-0.30})\times 10^{-14}$ erg cm-2 s-1.
We caution that all three Chandra DD observations (PIs: Li, O’Connor, and Eyles-Ferris) suffer from warm ACIS focal plane temperature’s [^5] https://cxc.cfa.harvard.edu/ciao/why/acis_fptemp.html with extreme episodes of heating and cooling throughout their exposures. Warm focal plane observations suffer from gain issues (gain under-corrected) and we caution against strong interpretations of the spectral analysis (especially due to the low 30-40 counts regime) without the proper gain corrections which are not updated yet for this observation period (Chandra Helpdesk [^6] https://cxc.harvard.edu/help/, private communication). This gain correction has likely impacted the spectral analysis presented by R. A. Eyles-Ferris et al. (2025) and EP250702a-arxiv.
@@ 3.1.1 Temporal Modeling
We modeled the 0.30.3-$10$ keV X-ray lightcurve (Swift/XRT and Chandra) with a powerlaw of the form $t^{-\alpha}$ using the emcee package (D. Foreman-Mackey et al., 2013). As the exact onset time of the transient is unclear (Table 2), we simplify this consideration by focusing on the start time of the afterglow $T_{0}$. This should not be confused with the onset time of the transient, which is set by the first detection by the Einstein Probe (H. Q. Cheng et al., 2025a; EP250702a-arxiv), despite this early detection preceding the peak of the X-ray and gamma-ray lightcurves. We use the combined Swift/XRT and Chandra X-ray lightcurve to determine the afterglow start time assuming it decays as a powerlaw. We performed multiple different fits to the data to explore the range of possible afterglow start times and temporal decay rates. We first allowed the afterglow start time $T_{0}$ to be a free parameter, and then performed an additional fit with $T_{0}$ fixed to the first high-energy trigger time (the discovery time) of GRB 250702D by Fermi (see Table 2).
@@
We performed these fits both with and without the inclusion of the initial XRT orbit at <<$0.55$ days after discovery (see Figure 14). We demonstrate that the initial XRT observation shows flaring activity (§3.1.2) and may bias the temporal analysis. Ignoring the initial flaring episode does not significantly change the inferred decay slope (see Figure 14), but does shift the inferred value of $T_{0}$ (Figure 1; right panel) towards earlier times. This demonstrates that the exact decay slope and afterglow start time depend strongly on the assumptions taken for the origin of the X-ray lightcurve.
@@
Using the trigger of GRB 250702D as $T_{0}$, we similarly modeled the NuSTAR 33-$79$ keV flux (see §3.2) across the three epochs (with mid-times of 1.58, 5.70, 9.98 d; Table 1) with a temporal powerlaw model. We derive a temporal index of $\alpha$$2.26\pm 0.15$. This is slightly steeper, but consistent at the $\sim$$2\sigma$ level, with the XRT fit using the same fixed start time (see Figure 14). We note that the NuSTAR data is limited to three epochs and the third epoch suffers from large error bars, which suggests that the slope derived from the combined Swift and Chandra lightcurve is more accurate. The first epoch of NuSTAR data also suffers from flaring (see §3.1.2), which can also explain its higher flux that can lead to a steeper inferred slope. The shallower slope measured from Swift and Chandra is also more consistent with the early Swift/BAT survey mode upper limits (see §2.1.1; Table 3). This is shown in Figure 5. While the NuSTAR extrapolation goes through two BAT detections of the prompt emission, this depends strongly on the assumed temporal slope and we emphasize the BAT detections are not included in the fit.
Figure 5: Left: Swift/BAT survey mode (1414-$195$ keV) on emission from GRB 250702B from the start of July 1, 2025 to the end of July 4, 2025. Black points show the Swift/BAT detections (1414-$195$ keV; Tables 5 and 4) from both GUANO and survey data. Gray downward triangles represent $5\sigma$ upper limits from Swift/BAT in the 1414-$195$ keV energy range. The trigger times reported by Fermi (red) and EP (blue) are shown as vertical lines (Table 2). The July 1 window in which EP reports an initial detection is shown as a blue shaded region. The time window of our initial NuSTAR observation (starting 1.3 d after the initial Fermi trigger) is shown as an orange shaded region. Right: Best-fit temporal powerlaw to the NuSTAR lightcurve in 33-$79$ keV. The $1\sigma$ error on the fit is shown as a shaded orange region. Swift/BAT upper limits are shown as gray triangles and detections as black circles. The start time $T_{0}$ is taken as the trigger time of GRB 250702D. The vertical red lines show the Fermi trigger times of GRBs 250702B, 250702C, and 250702E (see Table 2).
@@ 3.1.4 Average Power Spectrum
We note that a significantly shorter variability timescale of $\sim$11 s was found in the gamma-ray data from Fermi/GBM (Neights2025). However, we caution that the low count rate of the NuSTAR data cannot probe 1 s variability timescales and we cannot determine whether the lack of identified X-ray variability on shorter timescales during our observation is a physical property of the source.
@@ 3.2.1 X-ray Spectra
In all fits we have frozen the Galactic hydrogen column density to $N_{\textrm{H,gal}}$$3.34\times 10^{21}$ cm-2 (Willingale2013). The data were modeled between 0.3100.3-10 keV for Swift and 3793-79 keV for NuSTAR. Using an absorbed powerlaw model, we find an intrinsic hydrogen column density $N_{\textrm{H,z}}$$(6.2\pm 0.8)\times 10^{22}$ cm-2 and X-ray photon index $\Gamma$$1.82\pm 0.03$ with Cstat/dof ==$1277/1390$. The spectral fit and fit residuals are shown in Figure 7. We do not find evidence for spectral features [^9] We note the presence of residuals at $\sim$ 8 keV in FPMB, which were determined to be noise (Brian Grefenstette, private communication) and not a real feature., such as emission or absorption lines or reflection features. This further supports an extragalactic scenario for GRB 250702B as it would be peculiar for an X-ray binary in outburst.
@@ 4 The Peculiar High-Energy Properties of GRB 250702B
Figure 10: K-corrected rest frame (0.30.3-$10$ keV) X-ray lightcurves of Swift long duration GRBs, ultra-long GRBs (GRBs 111209A, 121027A, 130925A; Levan2014; P. A. Evans et al., 2014a), and relativistic jetted TDEs (J. S. Bloom et al., 2011; S. B. Cenko et al., 2012; G. C. Brown et al., 2015; Pasham2023) compared to GRB 250702B for a few different $T_{0}$. For the July 2nd start time (purple), we have used the trigger time of GRB 250702D as $T_{0}$ (Table 2). For the July 1 start time (blue), we use the beginning of the first EP detection window (Hui Sun, private communication).
@@ 4.1 Ultra-long Prompt Emission
While these properties are atypical for GRBs, they have likewise never been observed from a relativistic TDE. While relativistic TDEs have been uncovered in binned gamma-ray data (S. B. Cenko et al., 2012; G. C. Brown et al., 2015), or 1000 s long image triggers (Sakamoto2011J1644; D. N. Burrows et al., 2011), they have never triggered high-energy monitors in the standard “rate” trigger method that detects typical GRBs, and they have definitely not done so multiple times in a 24 hour window. As near all-sky gamma-ray monitors have been operating for decades, multiple gamma-ray triggers would be more difficult to miss in the past than early soft X-ray emission that could only have been found by the Einstein Probe which launched only in the last 2 years (Yuan2025). A detailed sub-threshold gamma-ray search in the day(s) preceding ultra-long GRBs is likely necessary to further investigate this possibility.
@@
An important clue can be obtained from the isotropic-equivalent gamma-ray energy $E_{\gamma,\textrm{iso}}$. The integrated energy obtained by Konus-Wind (D. Frederiks et al., 2025) is $E_{\gamma,\textrm{iso}}$$(1.4^{+0.4}_{-0.2})\times 10^{54}$ erg in the 11-$10,000$ keV energy range (see also Neights2025). This is at the high end of the isotropic-equivalent energies found in classical GRBs and ultra-long GRBs (see Figure 11). This is also substantially higher than the energy released ($\approx$$10^{51}$ erg; 11-$10,000$ keV) by Sw J1644+57 during its initial Swift/BAT detection (GRB 110328A; Sakamoto2011J1644). The individual energetics of the “DBE” events yield energies a factor of $10\times$ higher (Neights2025). However, Sw J1644+57 triggered BAT multiple times over the next two days before BAT triggers were disabled for the source (D. N. Burrows et al., 2011). We estimate the gamma-ray energy during this peak phase by adopting the average fluence reported by D. N. Burrows et al. (2011) over this 2 d window and adopting the spectral shape (peak energy $\sim$$70$ keV) reported by Levan2011. This yields $E_{\gamma,\textrm{iso}}\approx$$(1.0\pm 0.2)\times 10^{53}$ erg in the 11-$10,000$ keV. This value is derived during the peak phase ($\sim$22 d) of Sw J1644+57’s emission (in both X-ray and gamma-rays), but Sw J1644+57 was detected by BAT both before and after this peak phase (D. N. Burrows et al., 2011). Therefore, our derived energy during this 2 d window is a lower limit to its full gamma-ray energy release. We show this lower limit as an orange shaded region in Figure 11 for comparison to the energetics of GRB 250702B and classical (short, long, and ultra-long) GRBs. Another notable difference to GRB 250702B is that the gamma-ray emission from Sw J1644+57 was substantially softer (peak energy $\sim$$70$ keV; Levan2011) than the >>1 MeV emission observed from GRB 250702B (Neights2025).
@@
In addition, the Fermi data revealed a critical piece of evidence in the minimum variability time. Neights2025 identified a very short second duration ($\lesssim$11 s; observer frame) variability timescale, which is a factor of 100100-$1,000$ smaller than observed in relativistic TDEs (J. S. Bloom et al., 2011; Pasham2023). Relativistic TDEs, and accreting black holes in general, are limited by the Schwarzschild light crossing time $t_{s}$$2R_{\textrm{g}}/c$, where $R_{\textrm{g}}$$GM/c^{2}$ is the gravitational radius, which depends on the black hole mass. As larger masses lead to longer predicted minimum variability timescales, if a massive black hole is invoked it would require a small $\lesssim 5\times 10^{4}M_{\odot}$ intermediate-mass black hole. It is also important to note that the light crossing timescale argument is only a lower limit to the minimum variability timescale, and the black hole mass determined through this argument for past relativistic TDEs is larger than the latest analysis based on their jet shutoff timescales (T. Eftekhari et al., 2024).
However, we note that relativistic TDEs have been compared to a smaller version of a blazar (J. S. Bloom et al., 2011), which have been observed to have variability on timescales less than their light crossing times (D. Giannios et al., 2009; BlazVarNature). The shorter variability timescale observed from blazars has been explained by the “jets in a jet” model (D. Giannios, 2013). Whether the same magnetic reconnection mechanism can explain the prompt emission of GRB 250702B is outside the scope of this paper, but we caution that this minijet scenario would mean that the short variability timescale does not directly constrain the black hole mass.
@@
In any case, this seconds long variability is typical of that observed in long duration GRBs, and would therefore most obviously favor an interpretation that invokes a stellar-mass progenitor such as in GRBs. As an additional argument, in the collapsar progenitor channel, the central engine for long GRBs is likely a stellar-mass black hole (Woosley1993). Assuming a $\sim 5M_{\odot}$ black hole, the predicted minimum variability timescale based on the Schwarzschild light crossing time is $5\times 10^{-5}$ s, which is significantly smaller (by orders of magnitude) than the minimum variability observed in GRBs (see, e.g., $^{-2}-1001$$0$ s; V. Z. Golkhou & N. R. Butler, 2014). Following this argument, if applied to massive black holes, it would imply the minimum variability from light crossing is significantly smaller than the likely true intrinsic variability that can be observed. We therefore consider the second duration minimum variability of GRB 250702B (Neights2025) as strong evidence for a stellar-mass progenitor (§5.2).